a race car accelerates uniformly from 18.5m/s to 46.1m/s in 2.47 seconds. Determine the acceleration of the car and the distance traveled



1
Expert's answer
2021-12-12T16:41:10-0500

"v=v_0+at\to a=(v-v_0)/t=(46.1-18.5)/2.47="


"=11.2\ (m/s^2)"


"s=(v^2-v_0^2)/(2a)=(46.1^2-18.5^2)/(2\cdot11.2)=79.6\ (m)"